# Benchmarking Universal Machine Learning Interatomic Potentials for Elastic Property Prediction

### v1

@@ 3.2 Systematic Error Analysis
Figure 6 presents the relative error distributions of the four uMLIPs with respect to DFT values across various elastic descriptors and the values in parentheses in this paragraph correspond to median relative errors. CHGNet exhibits pronounced systematic deviations across most properties. In bulk modulus predictions, it shows a median error of $2.61\%$, indicating a tendency toward underestimation, whereas MACE and SevenNet display slight overestimations ($4.16\%$ and $2.88\%$, respectively), and MatterSim remains close to zero bias ($1.98\%$). For the shear and Young’s moduli, CHGNet strongly underestimates both ($48.02\%$ and $44.20\%$), in sharp contrast to the overestimations observed for MACE ($13.83\%$ and $12.43\%$) and SevenNet ($9.79\%$ and $8.89\%$), while MatterSim again yields nearly symmetric distributions ($2.12\%$ and $2.24\%$). For Poisson’s ratio, CHGNet systematically overestimates ($27.25\%$), opposite to the mild underestimations of MACE ($4.35\%$) and SevenNet ($3.40\%$), whereas MatterSim remains almost unbiased ($0.70\%$). The bulk/shear ratio further highlights CHGNet’s strong positive bias ($77.05\%$), while the other models show values close to zero. CHGNet also shows especially high variability in anisotropy metrics, suggesting instability in capturing complex anisotropic behavior. For Cauchy pressure, CHGNet exhibits a systematic positive bias, whereas the others lean toward negative deviations. The large deviations in the predicted anisotropy and Cauchy pressure mainly reflect the high sensitivity of these quantities to small differences among elastic constants. In this work, most materials exhibit a relatively small degree of elastic anisotropy, with a DFT average of 1.97. The Cauchy pressure, being a difference quantity defined as $(C_{12}-C_{44})$, has a DFT average value of 17.9 GPa, which is much smaller than the bulk and Young’s moduli that are generally on the order of 100 GPa. Consequently, even moderate relative errors in the stiffness components can lead to large percentage deviations in these derived quantities, indicating that further development of uMLIPs to improve their accuracy in elastic property predictions is essential.Finally, in Debye temperature predictions, CHGNet again underestimates ($25.89\%$), while MACE ($6.55\%$) and SevenNet ($4.89\%$) perform closer to DFT, and MatterSim achieves the most balanced performance ($0.69\%$). Overall, CHGNet displays consistent systematic biases across multiple properties, whereas MACE, MatterSim, and SevenNet yield more symmetric, near-zero error distributions, reflecting higher robustness.
@@
Figure 6: (a)-(h) Distribution of relative errors (%) for CHGNet, MACE, MatterSim, and SevenNet compared with DFT reference values across eight elastic properties. Each boxplot shows the median (blue dashed line), mean (red dashed line), interquartile range (colored box), and overall spread of errors (short lines indicating the minimum and maximum values), with outliers omitted for clarity. The long dashed line marks zero error.
@@
To provide a clearer comparison of overall performance, Figure 7 summarizes the MAPE values of different models across all elastic properties. It is evident that CHGNet systematically yields the highest error levels, with an average MAPE of 71.8%, underscoring its structural deficiencies in elastic property prediction. In contrast, SevenNet consistently achieves the lowest error, with an average MAPE of only 27.53%, highlighting its superior overall accuracy. Further analysis reveals that differences among models are relatively small for bulk modulus and Young’s modulus, whereas much larger discrepancies arise in shear modulus, the bulk/shear ratio, and Cauchy pressure—properties closely linked to mechanical stability and anisotropy. Particularly noteworthy is that CHGNet’s MAPE exceeds 90% for these metrics, reflecting structural limitations in capturing the complex couplings within elastic tensors. While MACE and MatterSim outperform CHGNet, their overall accuracy remains inferior to SevenNet, reinforcing the conclusion that SevenNet exhibits stronger generalization capability in modeling the nonlinear interdependencies among elastic properties.
@@
4 Discussion
4.1 Implications for Materials Design Applications
@@
The benchmark results discussed above provide clear guidance for selecting suitable uMLIPs according to specific application requirements. For tasks requiring highly accurate predictions of elastic properties, SevenNet should be prioritized; although its computational cost is somewhat higher, it offers more reliable performance. For high-throughput screening workflows, MACE and MatterSim strike a favorable balance between accuracy and efficiency, making them better suited for large-scale applications. While CHGNet shows comparatively weaker overall performance, it remains a viable option for simulations involving magnetic systems, where its specialized capabilities can be advantageous.
@@
Systematic bias patterns observed across all models warrant careful consideration in practical applications. In particular, consistent tendencies toward underestimation or overestimation of elastic moduli highlight the need for bias-correction strategies. For quantitative materials design, it is recommended that final results be validated against high-accuracy DFT calculations to ensure reliability.
4.2 Fundamental Limitations and Future Directions
@@
The current limitations of uMLIPs primarily stem from training datasets that are biased toward equilibrium configurations and lack adequate sampling of strained states, which are crucial for accurate elastic property predictions. Future developments should focus on systematically incorporating deformed structures into training datasets, for instance through active learning strategies that aim to improve mechanical property accuracy [45]. In addition, model fine-tuning[46] has emerged as a cost-effective optimization strategy and has already been widely adopted in other domains. For elastic property predictions within specific chemical spaces, constructing domain-specific fine-tuning [28] datasets and adapting pretrained models accordingly could effectively mitigate systematic biases in those regions.
@@
5 Conclusions
Our benchmark study establishes the first systematic evaluation framework for applying uMLIPs to elastic property prediction, validated across nearly 11,000 crystalline materials. The results demonstrate clear differences in model suitability: SevenNet delivers the highest overall accuracy, MatterSim and MACE achieve a favorable balance between accuracy and computational efficiency, while CHGNet, constrained by its architectural design, performs relatively less effectively. These findings provide not only evidence-based guidance for model selection in mechanical property calculations but also underscore the importance of tailoring model choice to specific application scenarios.
Comprehensive analyses of systematic biases further reveal common limitations among current uMLIPs, including the consistent under- or overestimation of elastic moduli and a training-data bias toward equilibrium configurations. Building on these insights, we identify several promising directions for future development, such as incorporating strained structures through active learning, implementing property-specific fine-tuning protocols, and establishing systematic error-correction schemes. We anticipate that such advances will further improve the reliability of uMLIPs for quantitative and high-throughput materials design, while also laying the groundwork for the next generation of universal interatomic potentials.
Acknowledgments
@@
Supporting Information
Figure S1: Processing time distribution comparison across machine learning interatomic potentials for elastic property calculations. Box plots show the statistical distribution of computational times, with boxes representing the interquartile range (25th-75th percentiles), horizontal lines indicating median and mean values, and whiskers extending to the furthest non-outlier data points. Outliers were filtered using z-score method (factor = 3.0) during preprocessing and are not displayed.

### v2

@@ III.2 Systematic Error Analysis
Fig. 6 presents the relative error distributions of the four uMLIPs with respect to DFT values across various elastic descriptors and the values in parentheses in this paragraph correspond to median relative errors. CHGNet exhibits pronounced systematic deviations across most properties. In bulk modulus predictions, it shows a median error of $2.61\%$, indicating a tendency toward underestimation, whereas MACE and SevenNet display slight overestimations ($4.16\%$ and $2.88\%$, respectively), and MatterSim remains close to zero bias ($1.98\%$). For the shear modulus and Young’s modulus, CHGNet strongly underestimates both ($48.02\%$ and $44.20\%$), in sharp contrast to the overestimations observed for MACE ($13.83\%$ and $12.43\%$) and SevenNet ($9.79\%$ and $8.89\%$), while MatterSim again yields nearly symmetric distributions ($2.12\%$ and $2.24\%$). For Poisson’s ratio, CHGNet systematically overestimates ($27.25\%$), opposite to the mild underestimations of MACE ($4.35\%$) and SevenNet ($3.40\%$), whereas MatterSim remains almost unbiased ($0.70\%$). The bulk/shear ratio further highlights CHGNet’s strong positive bias ($77.05\%$), while the other models show values close to zero. CHGNet also shows especially high variability in anisotropy metrics, suggesting instability in capturing complex anisotropic behavior. For Cauchy pressure, CHGNet exhibits a systematic positive bias, whereas the others lean toward negative deviations. The large deviations in the predicted anisotropy and Cauchy pressure mainly reflect the high sensitivity of these quantities to small differences among elastic constants. In this work, most materials exhibit a relatively small degree of elastic anisotropy, with a DFT average of 1.97. The Cauchy pressure, being a difference quantity defined as $(C_{12}-C_{44})$, has a DFT average value of 17.9 GPa, which is much smaller than the bulk and Young’s modulus that are generally on the order of 100 GPa. Consequently, even moderate relative errors in the stiffness components can lead to large percentage deviations in these derived quantities, indicating that further development of uMLIPs to improve their accuracy in elastic property predictions is essential. Finally, in Debye temperature predictions, CHGNet again underestimates ($25.89\%$), while MACE ($6.55\%$) and SevenNet ($4.89\%$) perform closer to DFT, and MatterSim achieves the most balanced performance ($0.69\%$). Overall, CHGNet displays consistent systematic biases across multiple properties, whereas MACE, MatterSim, and SevenNet yield more symmetric, near-zero error distributions, reflecting higher robustness.
@@
Figure 6: (a)–(h) Distribution of relative errors (%) for CHGNet, MACE, MatterSim, and SevenNet compared with DFT reference values across eight elastic properties. Each boxplot shows the median (blue dashed line), mean (red dashed line), interquartile range (colored box), and the overall error range (short lines marking the extrema), with outliers omitted for clarity. The horizontal dashed line denotes zero error.
@@
To provide a clearer comparison of overall performance, Fig. 7 summarizes the MAPE values of different models across all elastic properties. It is evident that CHGNet systematically yields the highest error levels, with an average MAPE of 71.8%, underscoring its structural deficiencies in elastic property prediction. In contrast, SevenNet consistently achieves the lowest error, with an average MAPE of only 27.53%, highlighting its superior overall accuracy. Further analysis reveals that differences among models are relatively small for bulk and Young’s modulus, whereas much larger discrepancies arise in the bulk/shear ratio, universal anisotropy index and Cauchy pressure—properties closely linked to mechanical stability and anisotropy. Particularly noteworthy is that CHGNet’s MAPE exceeds 90% for these metrics, reflecting structural limitations in capturing the complex couplings within elastic tensors. While MACE and MatterSim outperform CHGNet, their overall accuracy remains inferior to SevenNet, reinforcing the conclusion that SevenNet exhibits stronger generalization capability in modeling the nonlinear interdependencies among elastic properties.
@@
III.3 Fundamental Limitations and the Impact of Fine-Tuning
The limited performance of current uMLIPs in elastic property prediction can be partially attributed to training datasets dominated by equilibrium configurations, which do not provide sufficient coverage of strained states that are important for learning accurate mechanical responses. Fine-tuning [Fine-Tuning] provides a practical route to mitigating such systematic biases. To evaluate this strategy, we constructed a targeted fine-tuning dataset comprising 185 materials with the largest baseline errors, together with DFT-computed energies of their deformed structures, thereby explicitly introducing non-equilibrium information into the training domain. The DFT calculation settings are detailed in the Supporting Information. Fine-tuning on this targeted dataset yields well-controlled training mean absolute errors (MAEs): CHGNet achieved energy/force/stress MAEs of 24.089 meV/atom, 26.831 meV/Å, and 1.960 meV/Å3; MACE obtained 6.17 meV/atom, 26.38 meV/Å, and 5.67 meV/Å3; MatterSim reached 17.89 meV/atom, 9.574 meV/Å, and 0.836 meV/Å3; and SevenNet yielded 1.692 meV/atom, 5.101 meV/Å, and 0.528 meV/Å3.
All four uMLIPs—CHGNet, MACE, MatterSim, and SevenNet—were fine-tuned on this dataset. The updated potentials were then reassessed on the same 185 materials. Fig. S2 and Fig. S3 display heatmaps of the MAPE before and after fine-tuning, while Fig. S4 and Fig. S5 present the corresponding distributions of relative errors. To quantify these effects, Table 1 summarizes the relative change in MAPE, defined as the difference between the fine-tuned MAPE and the original MAPE, normalized by the original value. CHGNet shows the largest improvement, with an average MAPE reduction of 23.2%, followed by MatterSim at 20.7% and SevenNet at 18.0%. In contrast, MACE’s average MAPE increases by 13.8%, indicating a less favorable response to the added deformation information. Pronounced improvements occur for the two dimensionless ratios—the Poisson’s ratio and the bulk-to-shear modulus ratio—whose MAPE reductions exceed 40% for CHGNet, MatterSim, and SevenNet. In comparison, the six dimensional elastic constants (e.g., bulk modulus, shear modulus, Young’s modulus) typically show more modest reductions around 10–20%, highlighting a differential sensitivity to the inclusion of strained configurations. Table 2 reports the corresponding interquartile ranges (IQR), showing that fine-tuning leads to IQR reductions in seven of the eight elastic properties for CHGNet, and in all eight properties for both MatterSim and SevenNet. By contrast, MACE shows IQR reduction in only two properties and increased dispersion in the remaining six. These results demonstrate a clear difference in fine-tuning consistency among the four uMLIPs. For CHGNet, MatterSim, and SevenNet, the simultaneous improvements in both MAPE and IQR across all eight mechanical quantities indicate a coherent propagation of the energy–force–stress refinements into uniformly improved elastic responses. This reflects a high degree of physical consistency, where the optimization of the fundamental quantities translates directly into better predictions of derived mechanical observables. In contrast, the inconsistent behavior of MACE—improving in only a subset of properties while degrading in others—suggests weaker coupling between its learned potential-energy surface and its strain-dependent responses, pointing to limited fine-tuning consistency and reduced robustness to strained configurations. Such behavior illustrates that fine-tuning not only reduces systematic bias for CHGNet, MatterSim, and SevenNet but also yields more stable and concentrated error distributions. The degradation in MACE’s performance may reflect overfitting or architectural sensitivities that warrant further investigation.
These findings confirm that targeted fine-tuning with non-equilibrium configurations provides an effective and physically consistent route for improving the mechanical predictive capability of CHGNet, MatterSim, and SevenNet, whereas MACE shows limited and inconsistent gains.
Table 1: Relative percentage change in the MAPE of eight elastic properties for the four uMLIPs after fine-tuning.
Property CHGNet (%) MACE (%) MatterSim (%) SevenNet (%)
Bulk Modulus $K$ 11.8 -5.2 -16.5 -2.7
Shear Modulus $G$ -5.3 -14.6 -15.5 -5.7
Young’s Modulus $E$ -9.2 -9.9 -18.2 -4.5
Poisson’s Ratio $\nu$ -56.6 18.4 -42.4 -48.5
Bulk/Shear Ratio -55.4 89.4 -39.8 -41.3
Universal Anisotropy Index -31.6 37.6 -11.8 -13.9
Cauchy Pressure -22.4 -1.0 -9.6 -18.6
Debye Temperature -17.1 -4.3 -11.3 -9.1
Average -23.2 13.8 -20.7 -18.0
Table 2: IQR of the relative error for eight elastic properties before and after fine-tuning. Arrows indicate whether the IQR decreases ($\downarrow$) or increases ($\uparrow$).
Property CHGNet (%) MACE (%) MatterSim (%) SevenNet (%)
Bulk Modulus $K$ $236.2\rightarrow 242.4~(\uparrow)$ $268.7\rightarrow 250.4~(\downarrow)$ $270.6\rightarrow 208.6~(\downarrow)$ $262.6\rightarrow 260.2~(\downarrow)$
[…]
Young’s Modulus $E$ $272.7\rightarrow 236.6~(\downarrow)$ $251.0\rightarrow 258.5~(\uparrow)$ $267.6\rightarrow 217.1~(\downarrow)$ $258.6\rightarrow 241.9~(\downarrow)$
Poisson’s Ratio $\nu$ $69.4\rightarrow 29.0~(\downarrow)$ $36.1\rightarrow 52.4~(\uparrow)$ $46.2\rightarrow 20.6~(\downarrow)$ $37.5\rightarrow 19.5~(\downarrow)$
Bulk/Shear Ratio $88.2\rightarrow 53.0~(\downarrow)$ $51.5\rightarrow 104.2~(\uparrow)$ $60.8\rightarrow 35.5~(\downarrow)$ $54.2\rightarrow 28.8~(\downarrow)$
[…]
Cauchy Pressure $174.1\rightarrow 134.0~(\downarrow)$ $80.1\rightarrow 147.4~(\uparrow)$ $138.2\rightarrow 125.6~(\downarrow)$ $145.4\rightarrow 116.0~(\downarrow)$
Debye Temperature $99.5\rightarrow 78.9~(\downarrow)$ $85.2\rightarrow 99.9~(\uparrow)$ $91.4\rightarrow 77.0~(\downarrow)$ $86.6\rightarrow 81.6~(\downarrow)$
IV Discussion
IV.1 Implications for Materials Design Applications
@@
The benchmark and fine-tuning results discussed above provide clear guidance for selecting suitable uMLIPs according to specific application requirements. For tasks requiring highly accurate predictions of elastic properties, SevenNet should be prioritized; although its computational cost is somewhat higher, it offers more reliable performance. For high-throughput screening workflows, MACE and MatterSim strike a favorable balance between accuracy and efficiency, making them better suited for large-scale applications. While CHGNet shows comparatively weaker overall performance, it remains a viable option for simulations involving magnetic systems, where its specialized capabilities can be advantageous.
@@
Systematic bias patterns observed across all models warrant careful consideration in practical applications. In particular, consistent tendencies toward underestimation or overestimation of elastic modulus highlight the need for bias-correction strategies. For quantitative materials design, it is recommended that final results be validated against high-accuracy DFT calculations to ensure reliability.
IV.2 Future Directions
@@
As demonstrated in our fine-tuning study, incorporating strained configurations into the training process can effectively improve the predictive accuracy of uMLIPs for elastic properties, underscoring the importance of dataset diversity. Future developments should focus on systematically incorporating deformed structures into training datasets, for instance through active learning strategies that aim to improve mechanical property accuracy [zhang2020dp]. For elastic property predictions within specific chemical spaces, constructing domain-specific fine-tuning [deng2025systematic] datasets and adapting pretrained models accordingly could effectively mitigate systematic biases in those regions.
@@
V Conclusions
Our benchmark study establishes the first systematic evaluation framework for applying uMLIPs to elastic property prediction, validated across nearly 11,000 crystalline materials. The results demonstrate clear differences in model suitability: SevenNet delivers the highest overall accuracy, MatterSim and MACE achieve a favorable balance between accuracy and computational efficiency, while CHGNet, constrained by its architectural design, performs relatively less effectively. These results provide quantitative, evidence-based guidance for selecting suitable uMLIPs in mechanical property calculations, ensuring that model performance is properly matched to application requirements.
Comprehensive analyses of systematic biases further reveal common limitations among current uMLIPs, including the consistent under- or overestimation of elastic modulus and a training-data bias toward equilibrium configurations. Our fine-tuning results, incorporating strained configurations into model training, show that even limited data augmentation can effectively mitigate such biases and enhance model robustness. CHGNet, MatterSim, and SevenNet show uniform improvements across nearly all mechanical quantities, indicating coherent and physically consistent fine-tuning behavior, whereas MACE displays smaller and more variable gains. Building on these insights, we identify several promising directions for future development, such as incorporating strained structures through active learning, implementing property-specific fine-tuning protocols, and establishing systematic error-correction schemes. We anticipate that such advances will further improve the reliability of uMLIPs for quantitative and high-throughput materials design, while also laying the groundwork for the next generation of universal interatomic potentials.
Acknowledgements.
@@
Appendix A Supporting Information
The Supporting Information provides DFT computational details together with supplementary benchmarking analyses for the four uMLIPs evaluated in this work.
DFT Computational Details
All deformed configurations were computed using density functional theory (DFT) [kohn1965self] within the Vienna Ab initio Simulation Package (VASP) [VASP, VASP2]. The Perdew–Burke–Ernzerhof (PBE) generalized gradient approximation [PBE] and projector augmented-wave (PAW) [PAW] pseudopotentials were employed. To ensure strict consistency with the standardized Materials Project workflow, we adopted the same computational parameters: a plane-wave cutoff energy of 520 eV, electronic convergence set to $10^{-5}$ eV, and a $\Gamma$-centered Monkhorst–Pack $k$-point mesh [MP] generated with 1000/(number of atoms per cell) $k$-point. Spin polarization was enabled for all materials. For each strained configuration, a single-point DFT calculation was carried out to obtain total energies, forces, and stresses for use in fine-tuning the uMLIPs.
Supplementary Figures
Fig. S1 compares the processing-time distributions of the four uMLIPs for elastic-property calculations. Figs. S2 and S3 compare the mean absolute percentage errors for eight elastic properties before and after fine-tuning. Figs. S4 and S5 present the corresponding relative-error distributions.
Figure S1: Processing time distribution comparison across uMLIPs for elastic-property calculations. Each boxplot shows the median (solid line), mean (dashed line), and interquartile range (colored box), with short lines marking the extrema of the non-outlier data. Outliers were filtered using a z-score method (factor = 3.0) during preprocessing and are not displayed.
Figure S2: MAPE heatmap for elastic properties predicted by CHGNet, MACE, MatterSim, and SevenNet relative to DFT reference values before fine-tuning, for the selected 185 materials with high prediction errors. The properties analyzed include bulk modulus, shear modulus, Young’s modulus, Poisson’s ratio, bulk/shear ratio, universal anisotropy index, Cauchy pressure, and Debye temperature. Darker colors indicate higher errors, with values annotated in each cell.
Figure S3: MAPE heatmap for elastic properties predicted by CHGNet, MACE, MatterSim, and SevenNet relative to DFT reference values after fine-tuning, for the selected 185 materials with high prediction errors. The properties analyzed include bulk modulus, shear modulus, Young’s modulus, Poisson’s ratio, bulk/shear ratio, universal anisotropy index, Cauchy pressure, and Debye temperature. Darker colors indicate higher errors, with values annotated in each cell.
Figure S4: (a)-(h) Distribution of relative errors (%) for CHGNet, MACE, MatterSim, and SevenNet compared with DFT reference values across eight elastic properties for the selected 185 materials with high prediction errors before fine-tuning. Each boxplot shows the median (blue dashed line), mean (red dashed line), interquartile range (colored box), and the overall error range (short lines marking the extrema), with outliers omitted for clarity. The horizontal dashed line denotes zero error.
Figure S5: (a)-(h) Distribution of relative errors (%) for CHGNet, MACE, MatterSim, and SevenNet compared with DFT reference values across eight elastic properties for the selected 185 materials with high prediction errors after fine-tuning. Each boxplot shows the median (blue dashed line), mean (red dashed line), interquartile range (colored box), and the overall error range (short lines marking the extrema), with outliers omitted for clarity. The horizontal dashed line denotes zero error.